OpenStudy (anonymous):

5. A bag contains 6 silver and 5 gold coins. Which probability distribution represents the number of silver coins in a random draw of 4 coins?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {6 \over 11} \times {5 \over 10} \times {4 \over 9} \times {3 \over 8} }\)

OpenStudy (zarkon):

is your sampling with or without replacement?